OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5)

OpenStudy (anonymous):

find slope y=mx+c

OpenStudy (anonymous):

Uhh?

OpenStudy (binary3i):

y = -3x + 5

OpenStudy (anonymous):

@binary3i explain

OpenStudy (binary3i):

m is the slope m = (y2 - y1)/(x2 - x1) bye this you can get the slope after that wright the equ y =mx + c now put y1 at y and x1 at x this will give you the c. and it is done.